OpenStudy (anonymous):

how do i solve e^x+9=4

OpenStudy (lgbasallote):

solve for x?

OpenStudy (anonymous):

\[\ln4 -9=x\]

hartnn (hartnn):

first subtract 9 from both sides, u get e^x=-5 take log on both sides, u get x ln e = ln(-5) but since ln of negative number is not defined, this equation has 'no solution'. understood this ?

OpenStudy (anonymous):

@hartnn I actually think he might have meant \[e^{x+9} = 4\]

hartnn (hartnn):

ohh!! @sirena94 is it ?

hartnn (hartnn):

yup, it is e^(x+9)=4 so ln4-9 is correct.

OpenStudy (anonymous):

Alrighty! Thanks.